# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, skipabove=10pt, skipbelow=10pt

### v1

@@ V-B Autonomous Quadrotor Navigation
Next, we demonstrate our method on a 6-dimensional nonlinear quadrotor system. The system obeys the following dynamics:
@@
where $a_{g}=9.81\text{ m/s}^{2}$ is the acceleration due to gravity, and acceleration $u_{T}$, roll $u_{\phi}$, and pitch $u_{\theta}$ are the controls $u=[u_{T},u_{\phi},u_{\theta}]$ of the system with ranges $u_{T}\in[7.81,11.81]\text{ m/s}^{2}$ and $u_{\phi},u_{\theta}\in[-17.18^{\circ},17.18^{\circ}]$. Zero mean Gaussian noise is added to the actions to simulate real-world actuation errors.
The environment is a $10\text{m}\times 10\text{m}\times 4\text{m}$ room containing 10 spherical obstacles of varying sizes. The BRT of the system is computed using the LevelSetToolbox [21]. The task of the quadrotor is to reach a goal sphere with radius $0.3m$ around the goal state $\mathbf{x}_{g}$ (shown by the green sphere in Fig. 5), starting from a random initial position $\mathbf{x}_{0}$ while avoiding collisions with the obstacles. The system must reach the goal location within a time horizon of $T=10\text{s}$; otherwise, the episode is considered timeout. If the system collides with an obstacle or the environment boundary, the episode is considered unsafe.
The cost function for the MPPI algorithm consists of the distance to the goal location, control effort penalty, and a safety penalty $P(\mathbf{x})$:
@@
where $Q=\text{diag}([20,20,20,0,0,0])$, $u_{range}=[2,0.3,0.3]$ and $u_{mean}=[9.81,0,0]$. $P(\mathbf{x})$ is 10 if the state is inside an obstacle or the BRT, corresponding to that controller’s cost penalty type. For all methods, $500$ hallucinations are simulated for a horizon of $0.3s$. Each method is tested starting from the same 100 randomly sampled initial positions over the environment with at least $5m$ distance from the goal location. The results are summarized in Table II.
In this case, the obstacle penalty method leads to a very poor success rate. It is partially due to the added disturbance in the system that makes it challenging to ensure safety. A BRT penalty improves the safety rate from 14% to 30%, but it still does not enforce safety, as shown in Fig. 5. By penalizing a decrease in safety in the cost function and further optimizing the MPPI solution to encourage safety, Shield MPPI decreases the number of collisions to 21. In contrast, safety is guaranteed by a LRF of the controls, resulting in a 100% safety rate. Nevertheless, a significant portion of the hallucinations end up inside obstacles and are unable to contribute to the control synthesis meaningfully. Therefore, methods with safety-filtered actions experience many trajectories that fail to reach the goal within the given task horizon despite maintaining safety, in Fig. 5 we show such timeout executions. By filtering the hallucinations, our method incorporates the rollouts that would otherwise be unmeaningful into the optimization. Doing so achieves the best success and safety rates among all methods while achieving a similar cost on successful trajectories compared to the others that guarantee safety.
TABLE II: Quadrotor simulation results.
@@
Obs. penalty 14 0 86 0.43 $\pm$ 0.05
BRT penalty 30 0 70 0.50 $\pm$ 0.07
Obs. pen. + LRF 68 32 0 1.08 $\pm$ 0.35
BRT pen. + LRF 65 35 0 1.02 $\pm$ 0.38
Shield MPPI 69 10 21 0.84 $\pm$ 0.18
DualGuard (Ours) 75 25 0 1.00
@@ VI Hardware Experiments - RC Car
We adapt our evaluation metrics for this hardware study to better reflect real-world, single-run applicability. Instead of batch statistics, we measure the CompTime for each method – the time taken to generate and evaluate potential samples within a multiple-lap run. This metric reflects how well-suited each technique is for real-time control. We also report the car’s average Speed over three laps to measure how aggressive the policy is. The RelCost metric remains consistent with the simulations, providing a normalized cost relative to the proposed approach for the methods that managed to maintain safety.
@@
The controllers were implemented using JAX [30] on a laptop equipped with an NVIDIA GeForce RTX 4060. We generate $1000$ parallel hallucinations (with $100$ time steps each) in a loop running at $50Hz$. Results are summarized in Table III, and trajectories for the first lap are shown in Fig 6.
First, we highlight the need for hard safety constraints as the methods that only rely on safety penalties fail to clear the top-left tight turn in the track as shown in Fig 6. Fine-tuning the cost function and MPPI parameters might allow unfiltered methods to complete laps. Still, we want to consider and compare methods that provably allow for safe executions.
@@
The proposed method leads to faster and more performant trajectories than the other safe baselines. A direct comparison with the baselines that also use an output LRF illustrates that the proposed safe hallucination step improved the quality of the samples as exemplified in Fig. LABEL:fig:exp_main_result(B)(C), leading to a better overall performance and a higher average speed. Also, the proposed method outperforms the Shield-MPPI baseline even after tuning its hyperparameters to the best of our capabilities so that it maintains safety without an excessive impact on performance.
The computation times are nearly identical across all baselines, as each method fundamentally involves calculating performant terms of the cost function and querying the obstacle set or BRT for safety-related penalties. The proposed method introduces an additional LRF step for each sample along hallucinated trajectories, resulting in a slight increase in computational time. Nevertheless, all methods, including the proposed one, operate well within the $20ms$ time budget, leaving ample time for the control loop to handle state estimation, communications, and actuation.
@@ VII Limitations and Future Work
While these contributions establish DualGuard-MPPI as a robust and scalable framework, certain limitations remain that warrant further exploration. The requirement for a precomputed BRT introduces considerable computational overhead during setup, which may hinder its implementability in rapidly changing environments. Additionally, the reliance on explicitly defined system dynamics may restrict the framework’s applicability to systems with highly complex or partially unknown models. Addressing these challenges—such as by leveraging online reachability methods to dynamically update BRTs [31, 32, 33] or using reachability methods for black-box systems [34, 35] could significantly expand the framework’s usability. In addition, we will explore the deployment of the proposed approach on other safety-critical robotics applications.

### v2

@@ V-B Autonomous Quadrotor Navigation
Next, we demonstrate our method on a 6-dimensional nonlinear quadrotor system. The system follows the dynamics:
@@
where $a_{g}=9.81~{}\text{m/s}^{2}$ is gravity, and the controls $u=[u_{T},u_{\phi},u_{\theta}]$ correspond to thrust, roll, and pitch angles. Their ranges are $u_{T}\in[7.81,11.81]~{}\text{m/s}^{2}$ and $u_{\phi},u_{\theta}\in[-17.18^{\circ},17.18^{\circ}]$. Zero-mean Gaussian noise is added to simulate actuation errors.
The environment is a $10\text{m}\times 10\text{m}\times 4\text{m}$ room containing 10 spherical obstacles of varying sizes. $V(\mathbf{x})$ and the associated BRT of the system are numerically computed using the LevelSetToolbox [25]. The task of the quadrotor is to reach a goal sphere with radius $0.3m$ around the goal state $\mathbf{x}_{g}$ (shown by the green sphere in Fig. 3), starting from a random initial position $\mathbf{x}_{0}$ while avoiding collisions with the obstacles. The system must reach the goal location within a time horizon of $T=10\text{s}$; otherwise, the episode is considered timeout. If the system collides with an obstacle or the environment boundary, the episode is considered unsafe.
The cost function consists of distance to the goal, control effort penalty, and a method dependent safety penalty $P(\mathbf{x})$:
@@
where $Q=\text{diag}([20,20,20,0,0,0])$, $u_{range}=[2,0.3,0.3]$ and $u_{mean}=[9.81,0,0]$. For all methods, $500$ rollouts are simulated for a horizon of $0.3s$. Each method is tested starting from the same 100 randomly sampled initial positions over the environment with at least $5m$ distance from the goal location. The results are summarized in Table II.
In this case, the obstacle penalty method leads to a poor success rate. It is partially due to the added disturbance in the system that makes it challenging to ensure safety. A BRT penalty improves the safety rate from 14% to 30%, but it still does not enforce safety, as shown in Fig. 3. By penalizing decreases in safety and the approximate repair step , Shield MPPI decreases the number of collisions to 21. In contrast, safety is guaranteed for baselines with a LRF. Nevertheless, a significant portion of the rollouts end up inside obstacles and are unable to contribute to the control synthesis meaningfully. Therefore, methods with safety-filtered actions experience many trajectories that fail to reach the goal within the given task horizon despite maintaining safety; in Fig. 3, we show such timeout executions. By filtering the rollouts, our method incorporates samples that would otherwise be unmeaningful into the optimization. Doing so achieves the best success and safety rates among all methods and achieves better cost on safe trajectories compared to the others that guarantee safety.
TABLE II: Quadrotor simulation results. Asterisks (*) indicate statistically significant differences ($p<0.05$) compared to our method.
@@
Obs. penalty 14 0 86 0.48 $\pm$ 0.06*
BRT penalty 30 0 70 0.60 $\pm$ 0.09*
Obs. pen. + LRF 68 32 0 1.22 $\pm$ 0.06*
BRT pen. + LRF 65 35 0 1.21 $\pm$ 0.06*
Shield MPPI 69 10 21 0.91 $\pm$ 0.05
DualGuard (Ours) 75 25 0 1.00 $\pm$ 0.05
@@ VI Hardware Experiments - RC Car
We adapt our evaluation metrics to better reflect real-world, single-run applicability. Instead of batch statistics, we measure the CompTime for each method – the time taken to generate and evaluate potential samples within a multiple-lap run. This metric reflects how well-suited each technique is for real-time control. We also report the car’s average Speed over three laps to measure how aggressive the policy is. The RelCost metric remains consistent with the simulations.
@@
The controllers were implemented using JAX [33] on a laptop equipped with an NVIDIA GeForce RTX 4060. We generate $1000$ parallel rollouts (with $100$ time steps each) in a loop running at $50Hz$. Results are summarized in Table III, and trajectories for the first lap are shown in Fig 4.
First, we highlight the need for hard safety constraints as methods that only rely on safety penalties fail to clear the top-left tight turn as shown in Fig 4. Fine-tuning the cost function and MPPI parameters might allow unfiltered methods to complete laps. Still, we want to consider and compare methods that provably allow for safe executions.
@@
DualGuard leads to faster and more performant trajectories than the other safe baselines. Comparison with the LRF baselines shows how the proposed safe rollout step improved the quality of the samples as seen in Fig. LABEL:fig:exp_main_result(B)(C), leading to better overall performance and higher average speed. Also, the proposed method outperforms Shield-MPPI even after tuning its hyperparameters to the best of our capabilities to maintain safety without an excessive impact on performance.
Computation times are nearly identical across baselines, as each method fundamentally involves calculating performant terms of the cost function and querying the obstacle set or BRT for safety-related penalties. DualGuard introduces an additional LRF step for each sample along the trajectories, resulting in an increase in computational time. Nevertheless, all methods operate well within the $20ms$ time budget, leaving ample time for the control loop to handle state estimation, communications, and actuation.
@@ VII Limitations and Future Work
While these contributions establish DualGuard-MPPI as a robust and scalable framework, certain limitations remain that warrant further exploration. The requirement for a precomputed BRT introduces considerable computational overhead during setup, which may hinder its implementability in rapidly changing environments. Additionally, the reliance on explicitly defined system dynamics may restrict the framework’s applicability to systems with highly complex or partially unknown models. Addressing these challenges—such as by leveraging online reachability methods to dynamically update BRTs [34, 35] or using reachability methods for black-box systems [36, 37] could significantly expand the framework’s usability. In addition, we will explore the deployment of the proposed approach on other safety-critical robotics applications.